# Neutron star evolution by combining discontinuous Galerkin and finite volume methods

### v1

@@ IV.3.2 DG+FV/FD hybrid (S) simulation:
Lastly, we look at the central density $\rho_{0,\mathrm{C}}(t)$ normalized to the initial value (left plot) and the Hamiltonian constraint’s L2 norm in the central box (right plot) in Fig. 11, which are the representative quantities we study for all the DG+FV/FD simulations, including for (U) and (B) described later. For the Hamiltonian constraint, we compare all the (S) simulations. For $\rho_{0,\mathrm{C}}(t)$, we exclude the lowest resolution pure DG run as we know it has large noise in $\rho_{0,\mathrm{C}}(t)$. It should be kept in mind when making these comparisons, that the bulk of the mesh for our DG+FV/FD scheme has a resolution that falls in between the $l=4$ and $l=5$ cases for the pure DG simulations in Sec. IV.3.1. Even though the central h-refinement sphere for DG+FV/FD uses $l=5$, the dimension of the central cube is much larger than the previous setup. However, as the DG scheme is expected to demonstrate exponential convergence when increasing the order of the basis polynomials, the choice of $n=8$ for the DG elements in this DG+FV/FD evolution should give us much more accurate results for the same h-refinement level. As was mentioned in Sec. IV.3.1, the $l=4$ level was kept in the discussion only as a reference for improvement of results with increasing resolution. We started getting reliable results only from the $l=5$ resolution onwards. Keeping all these in mind, we expect the DG+FV/FD case to be the hybrid counterpart closest to the pure DG scheme of $l=5$ resolution. As can be seen from the left plot for $\rho_{0,\mathrm{C}}(t)$ in Fig. 11, the central density for DG+FV/FD essential settles down immediately into systematic oscillations compared to the pure DG runs. Additionally, the oscillation amplitude is much smaller. This is because the FV/FD at the surface of the star eliminates the Gibbs oscillations that plague the pure DG simulations at early times, when the perturbations from the numerical errors kick in. This is also reflected in the Hamiltonian constraint shown on the right of Fig. 11, which settles down much faster for the DG+FV/FD case. In fact, the results settle to a steady state even faster than the $l=6$ pure DG run, and to an even lower value. The comparable case of $l=5$ still remains at a level much above the DG+FV/FD case. Hence, we can see that the DG+FV/FD simulation performs significantly better than the equivalent pure DG simulation.
@@ IV.4 Perturbed configuration (P)
The remaining setup, apart from the pressure perturbation, remains exactly the same as the pure DG (S) simulation in Sec. IV.3.1. To study the effect of different resolutions, we again look at the central conserved density $D_{C}$ plot for the different refinement levels of $l=4,5$ and $6$, shown in Fig. 12. In this case, as there is an actual relatively stronger perturbation present, the oscillations in the central density are much more pronounced than in (S) where the oscillations were simply the result of much weaker perturbations form numerical noise. Again, we see that the $l=6$ case displays the expected behavior the best, while the lowest resolution of $l=4$ is once again swamped by Gibbs oscillations and numerical noise to the extent that we are unable to observe the characteristic oscillation we expect in this case. We again look at the Fourier transform of the central density in Fig.(6). Although the perturbation is significant, the deviation in the physical attributes such as mass and central density are within a margin such that characteristic spectrum is still in agreement with the (S) configuration as described in [121]. Hence, we still see the principal mode to be in perfect agreement with that predicted by theory. In fact, it is more pronounced as the oscillation is larger than (S). Once again, the higher modes are swamped by numerical errors.
@@ IV.5 Unstable configuration (U)
This case is significantly more challenging than the previous ones, as the matter fields are much more dynamic and there are shocks, which are most visible in the $\epsilon$ field [121, 47]. Even though the matter fields such as $\rho_{0}$ and $\epsilon$ were not smooth at the star surface in (S) and (P), the non-smooth feature remained relatively localized to its initial position as the configuration did not change drastically. The only motion came from the comparatively much gentler oscillations arising from either numerical or the pressure perturbation. However, the star (U) expands and contracts violently as it gets displaced from the initial configuration. It shoots past its stable configuration and goes back and forth between relatively higher and lower central density configurations. For handling these violent motions and shocks in the matter fields, the DG method alone was not suitable. Even though the system did evolve without crashing with DG alone, the dissipation caused by the filters and limiters was so significant, that the results, at affordable resolutions, were not close to the well known expected behavior as stated in [121, 122]. Hence, to handle this case we had to fall back to the FV/FD method, which provides robust HRSC capabilities.
@@
IV.5.1 Varying p-refinement for pure FV/FD (U) simulations:
[…]
For the case of p-refinement, we vary the number of points in each element in each direction to be $n=12,18,28$, or $42$, with $5$ elements across the star. Since we increase the number of points from the pure DG (S) and (P) setups and also made the central cube smaller, we lower the h-refinement level for these runs to $l=3$, so that the elements are larger, resulting in $6656$ total elements. The parameters for the limiters (see Sec. III.5.2) are the same as in Sec. IV.3.2. When converting from conserved to primitive variables, the procedure in Sec. III.6 is relaxed, and for this case we allow negative $\epsilon$ values by setting $\epsilon_{\mathrm{floor}}=-0.1$. Because of this, $\epsilon$ dips below 00 for short periods of time in places where a shock forms when matter falls back on to the star during contraction due to numerical errors in the lower resolutions $n=12$ and $n=18$. Yet, for the higher resolutions of $n=28$ and $n=42$, $\epsilon$ remains positive. We do not apply filters to any of the variables here, as we are evolving with pure FV/FD and our WENOZ2 scheme is capable of handling any non-smooth features. However, here we set $w=1$ in Eq. (III.2), thus dropping extrapolation near the boundary. At the boundary, when interpolation is needed for data exchange between neighbors, we use linear interpolation. One point away from element boundaries, we also add a second order dissipation operator to the GH system, by using the standard centered second derivative stencil there. The dissipation factor is set to $\sigma=0.1$. These extra modifications were done out of an abundance of caution. However, later studies have shown that they are not needed to keep the system stable. Consequently, they were not used in the more sophisticated hybrid simulations discussed below.
@@ IV.5.1 Varying p-refinement for pure FV/FD (U) simulations:
For the reasons discussed below, we shift the star away from the center of the mesh for these (U) simulations. Due to h-refinement, the corners of eight elements meet at the mesh center. If the star is centered on the mesh, sharp features in the density form at the center in our simulations. Similar sharp features have been observed for this unstable star system in [123, 83]. However, simulations with conventional FV schemes as in [121] do not show such sharp features. As explained in [83], we believe that these features cannot be easily resolved by conventional FV methods, and are therefore not shown in much of the literature. Hence, to obtain results as close as possible to the standard 2D or 3D simulations for this setup [121, 128, 47], and for ease of comparison and vetting of results, we have decided to move the mesh so that the star center (located at the $(0,0,0)$) is at the center of an element. This suppresses sharp features in our simulations, as our FV/FD scheme acts like any standard FV method at element centers. For the simulations discussed below, the mesh is centered on the point $(1,1,1)$. Even with this setup there still remain some oscillations in the central density $\rho_{0,\mathrm{C}}$. However, the overall nature of our results agrees with standard results obtained with FV methods in 3 dimensions. Additionally, the variation of $\rho_{0,\mathrm{C}}$ with resolution in left plot of Fig. (13) is consistent with second order convergence.
[…]
@@
IV.5.2 Varying h-refinement for pure FV/FD (U) simulations:
@@
Figure 14: Results for central h-refinements of $l=3,4$ and $5$ with $n=16$ for the pure FV/FD simulation of (U). Left: rest mass density $\rho_{0,\mathrm{C}}$ (normalized to the initial value); center: rest mass in central cube (normalized to the initial value); right: L2 norm of Hamiltonian constraint in central cube.
The radius of the most highly refined region is chosen to be $30.5$ to ensure that the bulk of the matter is contained inside this region. The refinement is dropped by one level after radii $30.5,61,122$ from the origin, with the radius of the entire mesh being $200$. The resultant initial number of elements across the star for this setup are stated in Table 2. The central region refinements are chosen to be $l=5,4,3$ for the different cases. The center of the mesh is shifted to $(1,1,1)$, $(0.5,0.5,0.5)$ and $(0.25,0.25,0.25)$ for the three different resolutions respectively, to always have the star center at the middle of an element, similar to Sec. IV.5.1. The FV/FD related setup for this run is the same as the treatment of the FV/FD elements in the DG+FV/FD run for (S) in Sec. IV.3.2. This means, we switch on the extrapolation near the element boundary for WENOZ2, use parabolic interpolation at the boundary, and limit dissipation to the interior of an element as in Eq. (34). The hydro setup also remains the same as in Sec. IV.3.2, and so does the GH system.
[…]
@@
We see the characteristic alternate peaks and troughs in the central density $\rho_{0,\mathrm{C}}$ with attenuating amplitude in the left plot, that is known to be the appropriate behavior of (U) setup when evolving with an EoS that allows for heating. In the right plot we see that the L2 norm of the Hamiltonian constraint also goes through the undulations as the matter flows in and out of the central box, while maintaining overall stability. Given that the lowest resolution shows stable and accurate behavior in the long term simulation, and the shorter runs show convergent behavior with increasing resolution for the various quantities, we can cautiously conclude that this setup for the simulations are handled well with Nmesh, and we expect that we will be able to see proper convergent behavior with longer runs and more different resolutions in our future work. Having said that, for these short duration runs, we find that $\rho_{0,\mathrm{C}}$ in Fig. 14 is consistent with convergence of order $2.55$ with resolution, while the Hamiltonian constraint L2 norm in Fig. 14 tends towards 2nd order convergence.
IV.5.3 DG+FV/FD (U) simulation:
Having confirmed the robustness and soundness of our new FV/FD method, we next evolved (U) with the hybrid DG+FV/FD method. For this case, the setup related to the mesh, WENOZ2, FV, FD, metric and hydro evolutions remain same as cases with h-refinement $l=3$ and $4$ and number of points $n=16$ in the previous section IV.5.2. However, we now incorporate our TEI exactly the same way we did for the DG+FV/FD simulation of (S) in Sec. IV.3.2. The associated treatments of all the aspects of the DG elements also remain the same as in Sec. IV.3.2. For these simulations, we apply a perturbation to $\rho_{0}$ of the form
@@ IV.5.3 DG+FV/FD (U) simulation:
We again check that our TEI is working properly by verifying that only the elements only near non-smooth regions switch to FV/FD. Figure 16 shows the density on the mesh wireframe, similar to Sec. IV.3.2, to check which parts have switched to FV/FD, while zooming in onto the central cube in the $xy$-plane. Unlike for (S), we cannot focus only on the central cube this time, as the matter flows out into the rest of the mesh over time. We pick four time frames to look at in this figure. The top left figure shows initial $t=0M_{\odot}$ setup, where we set every element to FV/FD. This is because the star surface has a much steeper rise in the matter fields, and so we simply set all elements to FV/FD initially to be cautious, letting the TEI set the appropriate elements to DG after eleven time steps.
@@
Figure 16: Part of the mesh in the $xy$-plane from (U) simulation with hybrid DG+FV/FD evolution method at four different times. Top left: $t=0M_{\odot}$ (initial state); top right: $t=136M_{\odot}$ (first minimum of $\rho_{0,\mathrm{C}}$); bottom left: $t=232M_{\odot}$ (next peak with $\rho_{0,\mathrm{C}}\sim 0.4\rho_{0,\mathrm{C}}(t=0)$); and bottom right $t=320M_{\odot}$ (second minimum of $\rho_{0,\mathrm{C}}$). FV/FD elements have denser equally spaced points, while DG has sparser unequally spaced points. We see that our hybrid scheme’s TEI successfully tracks the low density (shown through the colormap) troubled region as the star expands and contracts.
@@
The next plot (top right) at $t=136M_{\odot}$ is close to when the central density of the star reaches the first minimum after the star expands, in the context of left plot of Fig. 15. Only a narrow troubled region at the boundary of the star uses FV/FD, which has matter density between $10^{-2}\rho_{0,\mathrm{C}}(t=0)$ and $10^{-4}\rho_{0,\mathrm{C}}(t=0)$. The rest of the mesh elements are evolved with DG. Next (bottom left) time shown is $t=232M_{\odot}$, which is close to the time when the star first fully recontracts, producing the central maximum of $\rho_{0,\mathrm{C}}(t)\sim 0.4\rho_{0,\mathrm{C}}(t=0)$ in left plot of Fig. 15. The shell of FV/FD elements also follows the star surface. However, now the star is starting to get larger, with the density range $10^{-2}\rho_{0,\mathrm{C}}(t=0)$ to $10^{-4}\rho_{0,\mathrm{C}}(t=0)$ spreading further from the bulk of the matter in the center. It should be kept in mind that even though we are looking only at the $\rho_{0}$ field here, there are sharp features, and even shocks that are particularly visible in the $\epsilon$ field, linked to these fast moving low density shells, which has also been observed in [121, 47]. The TEI picks up on them and switch from DG to FV/FD as necessary. Hence, a thicker band of elements around the star switches to FV/FD. The center has also adopted FV/FD as there are oscillations at the center due to matter falling into it from all sides when the star contracts. Similar oscillations have been also observed in [83], when dissipation is low enough compared to more traditional schemes. Lastly, we show time $t=320M_{\odot}$ in the bottom right plot, corresponding to the second minimum in the left plot of Fig. 15 when the star has once again expanded. Here we see that the troubled region has spread considerably and most of the field of view in the plot is covered by FV/FD elements, while DG is used in the smooth interior of the star. Of course, if we zoom out, most of the rest of the mesh is still largely evolved with DG. From these plots, we can determine that our TEI functions properly, and can track the troubled region appropriately even in a case like (U) where the star is undergoing much more violent motions through expansion and contraction than the considerably benign case (S) of Sec. IV.3.2.
@@
We also check the behavior of $\rho_{0,\mathrm{C}}(t)$ and L2 norm of Hamiltonian constraint in the central cube for the DG+FV/FD runs with two different h-refinement levels, and compare them to the same quantities from our FV/FD runs with corresponding resolutions, as shown in Fig. 15 and Fig. 18. The lowest resolution runs provide comparison for longer time, as they are computationally cheapest. We see that the results for $\rho_{0,\mathrm{C}}(t)$ with DG+FV/FD runs match reasonably well with their same resolution pure FV/FD counterparts, considering the time offsets owing to difference in nature of the initial perturbations. The Hamiltonian contraint is lower for the same resolution for the DG+FV/FD simulation than the pure FV/FD simulation. Additionally, the Hamiltonian constraint improve from $l=3$ to $l=4$ resolution when considering only the DG+FV/FD curves, while $\rho_{0,\mathrm{C}}(t)$ shows similar behavior. These give us confirmation that the DG+FV/FD evolutions are performing well.
@@
Figure 17: Portion of the mesh in the $xy$-plane from the (U) simulation with the hybrid DG+FV/FD method with different central h-refinmenet levels. Top Left: $t=132M_{\odot}$ for $l=3$ and top right: $t=148M_{\odot}$ for $l=4$; bottom Left: $t=232M_{\odot}$ for $l=3$ and bottom right: $t=228M_{\odot}$ for $l=4$. Both show $\rho_{0}$ close to the time when $\rho_{0,\mathrm{C}}$ reaches its first minimum and the following maximum. We see that the width of the region evolved with FV/FD decreases with increasing h-refinement, as the troubled region is concentrated in the non-smooth region near the star surface.

### v2

@@ IV.3.2 DG+FV/FD hybrid (S) simulation
Lastly, we look at the central density $\rho_{0,\mathrm{C}}(t)$ normalized to the initial value (left plot) and the Hamiltonian constraint’s L2 norm in the central box (right plot) in Fig. 11, which are the representative quantities we study for all the DG+FV/FD simulations, including for (U) and (B) described later. For the Hamiltonian constraint, we compare all the (S) simulations. For $\rho_{0,\mathrm{C}}(t)$, we exclude the lowest resolution pure DG run as we know it has large noise in $\rho_{0,\mathrm{C}}(t)$. It should be kept in mind when making these comparisons that the bulk of the mesh for our DG+FV/FD scheme has a resolution that falls in between the $l=4$ and $l=5$ cases for the pure DG simulations in Sec. IV.3.1. Even though the central h-refinement sphere for DG+FV/FD uses $l=5$, the dimension of the central cube is much larger than the previous setup. However, as the DG scheme is expected to demonstrate exponential convergence when increasing the order of the basis polynomials, the choice of $n=8$ for the DG elements in this DG+FV/FD evolution should give us much more accurate results for the same h-refinement level. As was mentioned in Sec. IV.3.1, the $l=4$ level was kept in the discussion only as a reference for improvement of results with increasing resolution. We started getting reliable results only from the $l=5$ resolution onwards. Keeping all these in mind, we expect the DG+FV/FD case to be the hybrid counterpart closest to the pure DG scheme of $l=5$ resolution. As can be seen from the left plot for $\rho_{0,\mathrm{C}}(t)$ in Fig. 11, the central density for DG+FV/FD essential settles down immediately into systematic oscillations compared to the pure DG runs. Additionally, the oscillation amplitude is much smaller. This is because the FV/FD at the surface of the star eliminates the Gibbs oscillations that plague the pure DG simulations at early times, when the perturbations from the numerical errors kick in. This better behavior of the central density oscillations are also reflected in the corresponding spectrum in Fig. 6. We see that the first three peaks for the (S) DG+FV/FD case are much more prominent than for the pure DG (S) simulation.
This improvement is also reflected in the Hamiltonian constraint shown on the right of Fig. 11, which settles down much faster for the DG+FV/FD case. In fact, the results settle to a steady state even faster than the $l=6$ pure DG run, and to an even lower value. The comparable case of $l=5$ still remains at a level much above the DG+FV/FD case. Hence, we can see that the DG+FV/FD simulation performs significantly better than the equivalent pure DG simulation.
@@ IV.4 Perturbed configuration (P)
The remaining setup, apart from the pressure perturbation, remains exactly the same as the pure DG (S) simulation in Sec. IV.3.1. To study the effect of different resolutions, we again look at the central conserved density $D_{C}$ plot for the different refinement levels of $l=4,5$, and $6$, shown in Fig. 12. In this case, as there is an actual relatively stronger perturbation present, the oscillations in the central density are much more pronounced than in (S) where the oscillations were simply the result of much weaker perturbations from numerical noise. Again, we see that the $l=6$ case displays the expected behavior the best, while the lowest resolution of $l=4$ is once again swamped by Gibbs oscillations and numerical noise to the extent that we are unable to observe the characteristic oscillation we expect in this case. We again look at the Fourier transform of the central density in Fig. (6) for the $l=6$ h-refinement case. The finest element dimension and grid spacing detail for this simulation is the same as mentioned for the corresponding $l=6$ pure DG (S) case in Sec. IV.3.1. Although the perturbation is significant, the deviation in the physical attributes such as mass and central density are within a margin such that characteristic spectrum is still in agreement with the (S) configuration as described in [123]. Hence, we still see the principal mode to be in perfect agreement with that predicted by theory. In fact, it is more pronounced as the oscillation is larger than (S). Once again, the higher modes are swamped by numerical errors.
@@ IV.5 Unstable configuration (U)
This case is significantly more challenging than the previous ones, as the matter fields are much more dynamic and there are shocks, which are most visible in the $\epsilon$ field [123, 47]. Even though the matter fields such as $\rho_{0}$ and $\epsilon$ were not smooth at the star surface in (S) and (P), the nonsmooth feature remained relatively localized to its initial position as the configuration did not change drastically. The only motion came from the comparatively much gentler oscillations arising from either numerical or the pressure perturbation. However, the star (U) expands and contracts violently as it gets displaced from the initial configuration. It shoots past its stable configuration and goes back and forth between relatively higher and lower central density configurations. For handling these violent motions and shocks in the matter fields, the DG method alone was not suitable. Even though the system did evolve without crashing with DG alone, the dissipation caused by the filters and limiters was so significant, that the results, at affordable resolutions, were not close to the well-known expected behavior as stated in [123, 124]. Hence, to handle this case we had to fall back to the FV/FD method, which provides robust HRSC capabilities.
@@
IV.5.1 Varying p-refinement for pure FV/FD (U) simulations
Figure 13: Results for p-refinements of $n=12,18,28$, and $42$ and h-refinememnt level of $l=3$ for the pure FV/FD simulation of (U) till the first peak when the star contracts. Left: central rest mass density $\rho_{0,\mathrm{C}}$(normalized to the initial value); center: rest mass $\int DdV$ in the central cube (normalized to the initial value); right: L2 norm of Hamiltonian constraint in the central cube. We see that the peak value of $\rho_{0,\mathrm{C}}$ is in good agreement with that in other simulations such as in [123]. The offset in time of the different resolutions is because the star gets displaced from the initial state at different times due to difference in perturbations from the different resolutions. The rest mass and the Hamiltonian constraint show convergent behavior.
For the case of p-refinement, we vary the number of points in each element in each direction to be $n=12,18,28$, or $42$, with five elements across the star. Since we increase the number of points from the pure DG (S) and (P) setups and also made the central cube smaller, we lower the h-refinement level for these runs to $l=3$, so that the elements are larger, resulting in $6656$ total elements. The parameters for the limiters (see Sec. III.5.2) are the same as in Sec. IV.3.2. When converting from conserved to primitive variables, the procedure in Sec. III.6 is relaxed, and for this case we allow negative $\epsilon$ values by setting $\epsilon_{\mathrm{floor}}=-0.1$. Because of this, $\epsilon$ dips below 0 for short periods of time in places where a shock forms when matter falls back onto the star during contraction due to numerical errors in the lower resolutions $n=12$ and $n=18$. Yet, for the higher resolutions of $n=28$ and $n=42$, $\epsilon$ remains positive. We do not apply filters to any of the variables here, as we are evolving with pure FV/FD and our WENOZ2 scheme is capable of handling any nonsmooth features. However, here we set $w=1$ in Eq. (III.2), thus dropping extrapolation near the boundary. At the boundary, when interpolation is needed for data exchange between neighbors, we use linear interpolation. One point away from element boundaries, we also add a second-order dissipation operator to the GH system, by using the standard centered second-derivative stencil there. The dissipation factor is set to $\sigma=0.1$. These extra modifications were done out of an abundance of caution. However, later studies have shown that they are not needed to keep the system stable. Consequently, they were not used in the more sophisticated hybrid simulations discussed below.
@@ IV.5.1 Varying p-refinement for pure FV/FD (U) simulations
For the reasons discussed below, we shift the star away from the center of the mesh for these (U) simulations. Due to h-refinement, the corners of eight elements meet at the mesh center. If the star is centered on the mesh, sharp features in the density form at the center in our simulations. Similar sharp features have been observed for this unstable star system in [125, 83]. However, simulations with conventional FV schemes as in [123] do not show such sharp features. As explained in [83], we believe that these features cannot be easily resolved by conventional FV methods, and are therefore not shown in much of the literature. Hence, to obtain results as close as possible to the standard 2D or 3D simulations for this setup [123, 130, 47], and for ease of comparison and vetting of results, we have decided to move the mesh so that the star center [located at the $(0,0,0)$] is at the center of an element. This suppresses sharp features in our simulations, as our FV/FD scheme acts like any standard FV method at element centers. For the simulations discussed below, the mesh is centered on the point $(1,1,1)$. Even with this setup there still remain some oscillations in the central density $\rho_{0,\mathrm{C}}$. However, the overall nature of our results agrees with standard results obtained with FV methods in three dimensions. Additionally, the variation of $\rho_{0,\mathrm{C}}$ with resolution in left plot of Fig. (13) is consistent with second-order convergence.
From the center plot in Fig. 13 we see that mass conservation improves with resolution. For the highest resolution case, the mass remains almost constant until about a time of $t=100M_{\odot}$. Until this point, the bulk of the matter is contained in the innermost cube, which is the best resolved region. However, since the star is expanding, the matter gradually spills over into the surrounding cubed sphere patches. Since linear interpolation takes place at the element boundary between an element in the central cubical patch and an adjacent element in the cubed sphere region, the mass does not remain exactly conserved, causing the drop we see around this time. Furthermore, the resolution also drops in the elements in the cubed sphere region, which also contributes to the mass loss we observe. Lastly, the limiters and atmosphere treatment we apply perform pointwise fixes, that also contribute to this violation. At around $t=160M_{\odot}$ the central higher density region, containing matter with density above $10^{-2}\rho_{0,\mathrm{C}}(t=0)$, stops expanding, and starts contracting again. However, a lower density shell of about $10^{-2}\rho_{0,\mathrm{C}}(t=0)$ to $10^{-5}\rho_{0,\mathrm{C}}(t=0)$ continues to expand, going further into the less resolved region, which causes the $\int DdV$ quantity to keep dropping. This shell stops expanding at around $t=192M_{\odot}$, which is when we see the second stabilization of the mass. The overall takeaway again is that the mass conservation improves drastically with increasing resolution. The error caused by the interpolation, and the other factors converge away with increasing resolution, providing reliable results as we go to reasonably good resolutions.
@@
IV.5.2 Varying h-refinement for pure FV/FD (U) simulations
@@
Figure 14: Results for central h-refinements of $l=3,4$, and $5$ with $n=16$ for the pure FV/FD simulation of (U). Left: rest mass density $\rho_{0,\mathrm{C}}$ (normalized to the initial value); center: rest mass in central cube (normalized to the initial value); right: L2 norm of Hamiltonian constraint in central cube.
The radius of the most highly refined region is chosen to be $30.5$ to ensure that the bulk of the matter is contained inside this region. The refinement is dropped by one level after radii $30.5,61,122$ from the origin, with the radius of the entire mesh being $200$. The resultant initial number of elements across the star for this setup are stated in Table 2. The central region refinements are chosen to be $l=3,4,5$ for the different cases. The center of the mesh is shifted to $(1,1,1)$, $(0.5,0.5,0.5)$, and $(0.25,0.25,0.25)$ for the three different resolutions respectively, to always have the star center at the middle of an element, similar to Sec. IV.5.1. The FV/FD related setup for this run is the same as the treatment of the FV/FD elements in the DG+FV/FD run for (S) in Sec. IV.3.2. This means we switch on the extrapolation near the element boundary for WENOZ2, use parabolic interpolation at the boundary, and limit dissipation to the interior of an element as in Eq. (34). The hydro setup also remains the same as in Sec. IV.3.2, and so does the GH system.
The behavior of the different quantities such as $\rho_{0,\mathrm{C}}$, $\int DdV$, and L2 norm of the Hamiltonian constraint in the central cube remain similar to that for varying p-refinement, as can be seen from the left, center, and right plots in Fig. 14 respectively. We see systematic improvements in all the quantities with increasing resolution. Unfortunately, due to high computational cost, we could only continue the highest resolution simulation to barely beyond the first expansion phase. This limits our capability to make strong conclusions, but we can still see that the results improve with resolution and are trending in a desirable fashion. The lowest resolution simulation was continued for much longer, the results for which can be seen in Fig. 15 by looking at the dashed line curves.
@@
We see the characteristic alternate peaks and troughs in the central density $\rho_{0,\mathrm{C}}$ with attenuating amplitude in the left plot, that is known to be the appropriate behavior of (U) setup when evolving with an EoS that allows for heating. In the right plot we see that the L2 norm of the Hamiltonian constraint also goes through the undulations as the matter flows in and out of the central box, while maintaining overall stability. Given that the lowest resolution shows stable and accurate behavior in the long term simulation, and the shorter runs show convergent behavior with increasing resolution for the various quantities, we can cautiously conclude that this setup for the simulations are handled well with Nmesh, and we expect that we will be able to see proper convergent behavior with longer runs and more different resolutions in our future work. Having said that, for these short duration runs, we find that $\rho_{0,\mathrm{C}}$ in Fig. 14 is consistent with convergence of order $2.55$ with resolution, while the Hamiltonian constraint L2 norm in Fig. 14 tends towards second-order convergence.
IV.5.3 DG+FV/FD (U) simulation
Having confirmed the robustness and soundness of our new FV/FD method, we next evolved (U) with the hybrid DG+FV/FD method. For this case, the setup related to the mesh, WENOZ2, FV, FD, metric and hydro evolutions remain same as cases with h-refinement $l=3$ and $4$ and number of points $n=16$ in the previous Sec. IV.5.2. However, we now incorporate our TEI exactly the same way we did for the DG+FV/FD simulation of (S) in Sec. IV.3.2. The associated treatments of all the aspects of the DG elements also remain the same as in Sec. IV.3.2. For these simulations, we apply a perturbation to $\rho_{0}$ of the form
@@ IV.5.3 DG+FV/FD (U) simulation
We again check that our TEI is working properly by verifying that only the elements near nonsmooth regions switch to FV/FD. Figure 16 shows the density on the mesh wireframe, similar to Sec. IV.3.2, to check which parts have switched to FV/FD, while zooming in on the central cube in the $xy$-plane. Unlike for (S), we cannot focus only on the central cube this time, as the matter flows out into the rest of the mesh over time. We pick four time frames to look at in this figure. The top left figure shows initial $t=0M_{\odot}$ setup, where we set every element to FV/FD. This is because the star surface has a much steeper rise in the matter fields, and so we simply set all elements to FV/FD initially to be cautious, letting the TEI set the appropriate elements to DG after 11 time steps.
@@
Figure 16: Part of the mesh in the $xy$-plane from (U) simulation with hybrid DG+FV/FD evolution method at four different times. Top left: $t=0M_{\odot}$ (initial state); top right: $t=136M_{\odot}$ (first minimum of $\rho_{0,\mathrm{C}}$); bottom left: $t=232M_{\odot}$ (next peak with $\rho_{0,\mathrm{C}}\sim 0.4\rho_{0,\mathrm{C}}(t=0)$); and bottom right $t=320M_{\odot}$ (second minimum of $\rho_{0,\mathrm{C}}$). FV/FD elements have denser equally spaced points, while DG has sparser unequally spaced points. We see that our hybrid scheme’s TEI successfully tracks the low-density (shown through the colormap) troubled region as the star expands and contracts.
@@
The next plot (top right) at $t=136M_{\odot}$ is close to when the central density of the star reaches the first minimum after the star expands, in the context of the left plot of Fig. 15. Only a narrow troubled region at the boundary of the star uses FV/FD, which has matter density between $10^{-2}\rho_{0,\mathrm{C}}(t=0)$ and $10^{-4}\rho_{0,\mathrm{C}}(t=0)$. The rest of the mesh elements are evolved with DG. The next (bottom left) time shown is $t=232M_{\odot}$, which is close to the time when the star first fully recontracts, producing the central maximum of $\rho_{0,\mathrm{C}}(t)\sim 0.4\rho_{0,\mathrm{C}}(t=0)$ in the left plot of Fig. 15. The shell of FV/FD elements also follows the star surface. However, now the star is starting to get larger, with the density range $10^{-2}\rho_{0,\mathrm{C}}(t=0)$ to $10^{-4}\rho_{0,\mathrm{C}}(t=0)$ spreading further from the bulk of the matter in the center. It should be kept in mind that even though we are looking only at the $\rho_{0}$ field here, there are sharp features, and even shocks that are particularly visible in the $\epsilon$ field, linked to these fast-moving low-density shells, which has also been observed in [123, 47]. The TEI picks up on them and switches from DG to FV/FD as necessary. Hence, a thicker band of elements around the star switches to FV/FD. The center has also adopted FV/FD as there are oscillations at the center due to matter falling into it from all sides when the star contracts. Similar oscillations have been also observed in [83], when dissipation is low enough compared to more traditional schemes. Lastly, we show time $t=320M_{\odot}$ in the bottom right plot, corresponding to the second minimum in the left plot of Fig. 15 when the star has once again expanded. Here we see that the troubled region has spread considerably and most of the field of view in the plot is covered by FV/FD elements, while DG is used in the smooth interior of the star. Of course, if we zoom out, most of the rest of the mesh is still largely evolved with DG. From these plots, we can determine that our TEI functions properly, and can track the troubled region appropriately even in a case like (U) where the star is undergoing much more violent motions through expansion and contraction than the considerably benign case (S) of Sec. IV.3.2.
@@
We also check the behavior of $\rho_{0,\mathrm{C}}(t)$ and L2 norm of Hamiltonian constraint in the central cube for the DG+FV/FD runs with two different h-refinement levels, and compare them to the same quantities from our FV/FD runs with corresponding resolutions, as shown in Figs. 15 and 18. The lowest resolution runs provide comparison for longer time, as they are computationally cheapest. We see that the results for $\rho_{0,\mathrm{C}}(t)$ with DG+FV/FD runs match reasonably well with their same resolution pure FV/FD counterparts, considering the time offsets owing to difference in nature of the initial perturbations. The Hamiltonian constraint is lower for the same resolution for the DG+FV/FD simulation than the pure FV/FD simulation. Additionally, the Hamiltonian constraint improves from $l=3$ to $l=4$ resolution when considering only the DG+FV/FD curves, while $\rho_{0,\mathrm{C}}(t)$ shows similar behavior. These give us confirmation that the DG+FV/FD evolutions are performing well.
@@
Figure 17: A portion of the mesh in the $xy$-plane from the (U) simulation with the hybrid DG+FV/FD method with different central h-refinmenet levels. Top left: $t=132M_{\odot}$ for $l=3$; top right: $t=148M_{\odot}$ for $l=4$; bottom left: $t=232M_{\odot}$ for $l=3$; and bottom right: $t=228M_{\odot}$ for $l=4$. Both show $\rho_{0}$ close to the time when $\rho_{0,\mathrm{C}}$ reaches its first minimum and the following maximum. We see that the width of the region evolved with FV/FD decreases with increasing h-refinement, as the troubled region is concentrated in the nonsmooth region near the star surface.